OpenStudy (anonymous):

Ms. Smith has 30 students in her class: 20 are boys and 10 are girls. Mr. Jones also has 30 students in his class: 15 are boys and 15 are girls. One student is selected from each class to be on the student council. What is the probability that 2 boys are selected? A. 1/6 B. 1/3 C. 1/2 D. 7/12

OpenStudy (anonymous):

what do you think the total possible combination is?

OpenStudy (anonymous):

I am not sure but I was thinking A or C

OpenStudy (anonymous):

well, let's look at this way. How many students are there in Ms. Smith class room?

OpenStudy (anonymous):

Ms Smith has 30 students

OpenStudy (anonymous):

good. How about Mr. Jones class?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

good. So if select one student form each class, and pair them up. How many pairs can you form?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

no no, how many PAIRS can be formed?

OpenStudy (anonymous):

Oh, 15 pairs

OpenStudy (anonymous):

no. it's 30 * 30 = 900 pairs possible. Why?

OpenStudy (anonymous):

Because they have 30 students from one class and 30 students from the other class?

OpenStudy (anonymous):

Good. Now how many boys are there in Ms. Smith class?

OpenStudy (anonymous):

20 boys

OpenStudy (anonymous):

good. What about the other one?

OpenStudy (anonymous):

15 boys

OpenStudy (anonymous):

so if you select each boy from each class, how many pairs of boys can be formed?

OpenStudy (anonymous):

There are 15 pairs?

OpenStudy (anonymous):

no, it's 20 * 15 = 300

OpenStudy (anonymous):

so the probability that a pair contains two boys is?

OpenStudy (anonymous):

Do you always multiple in probability?

OpenStudy (anonymous):

no, you don't always multiply.

OpenStudy (anonymous):

1 pair

OpenStudy (anonymous):

???

OpenStudy (anonymous):

P = 300/900 yes?

OpenStudy (anonymous):

So, 300 times 2=900 because the 300 part is from the two class and the 2 is from the two boys.

OpenStudy (anonymous):

two classes*

OpenStudy (anonymous):

300 * 2 isn't 900. answer is 300/900 = 1/3 so B

OpenStudy (anonymous):

I feel like you still haven't fully understood why we did what we did.

OpenStudy (anonymous):

Yeah, a little bit. I hardly learned about probability.

OpenStudy (anonymous):

But, I truly thank you for using your time helping me step by step.

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

huhm... that's probably why. Probability isn't easy for middle high schoolers. Not because it's hard but because probabilities requires thinking logically, which is too advanced for a 9th grader.

OpenStudy (anonymous):

Pretty suck for me. But, I am willing to learn! :) Thank you

OpenStudy (anonymous):

it takes times and practices. As long as you're willing to learn, you'll get better :)